OpenStudy (anonymous):

Part 1: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in standard form to represent the total rent that same has to pay for renting the boat for x days. Part 2: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labes on the axes and the intervals. Please help!! D:

OpenStudy (anonymous):

principal amount+days*priceperday

OpenStudy (anonymous):

How would that look as a normal equation?

OpenStudy (anonymous):

im trying to figure it out :S

OpenStudy (anonymous):

Same. :/

OpenStudy (anonymous):

ok i think

OpenStudy (anonymous):

the price per day is not the same

OpenStudy (anonymous):

i did y=225-2x y=480-5x and solved for y got y=55

OpenStudy (anonymous):

So, what would be the answer for Part 1?

OpenStudy (anonymous):

55 p[rincipal + 85 per day

OpenStudy (anonymous):

price of rental=85x+55

OpenStudy (anonymous):

Ah, okay. Thanks!

OpenStudy (anonymous):

it works for both cases

OpenStudy (anonymous):

Gotcha. So, do you think you know the answer to Part 2?

OpenStudy (anonymous):

therefore they cahrge a free of 55 bucks + 85 days per day of rental

OpenStudy (anonymous):

easy, the graphs starts at x=0, and y=55, because thats the minimum they will eb cahrged at 0 days

OpenStudy (anonymous):

Also, {x + 2 y = 225, x + 5 y = 480} {x = 55, y = 85} x is how much to rent the boat and y is the daily rate.

OpenStudy (anonymous):



OpenStudy (anonymous):

the coordinates are as follows

OpenStudy (anonymous):

0,55 1,(55+85)

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

So, how should I answer Part 2?

OpenStudy (anonymous):

Hello? .3.

OpenStudy (anonymous):

A plot is attached. The plot is of the line y = 85x + 55 and the points { {2, 225}, {5, 480} }